# Wilsonian Renormalization of Neural Network Gaussian Processes

### v2

@@ III Renormalizing Gaussian features
$I_{0}[\Phi_{>}]+$ $I_{\mathrm{int}}[\Phi_{<},\Phi_{>}]$ (26)
$\coloneqq$ $\,\frac{1}{2\sigma^{2}}\sum_{m;kk^{\prime}>\kappa}(f_{mk}-y_{k})(f_{mk^{\prime}}-y_{k^{\prime}})\varphi_{k}\varphi_{k^{\prime}}$
@@
We can then express this result in terms of a covariance matrix of replica indices by Taylor expanding the logarithm in the coefficient and using the cyclicity of the trace to obtain
$\frac{1}{2}\mathrm{tr}\ln\left(\mathbb{1}_{N_{>}\times N_{>}}+\right.$ $\left.F^{\top}F/\sigma^{2}\right)$ (27)
$=\frac{1}{2}\mathrm{tr}\left(\frac{FF^{T}}{2\sigma^{2}}-\frac{(FF^{\top})^{2}}{4\sigma^{2}}+\ldots\right)$
@@
where $C\coloneqq FF^{\top}$ is an $M\!\times\!M$ covariance matrix with elements
$[C]_{mn}=(f_{mk}-y_{k})(f_{nk}-y_{k})\leavevmode\nobreak\ .$ (28)
@@
The argument of the exponential can similarly be written in terms of this matrix by applying the Woodbury matrix identity, whereupon
$F\left(\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\right)^{-1}\!F^{\top}=\left(C^{-1}+\sigma^{-2}\right)^{-1}\leavevmode\nobreak\ .$ (29)
@@
Putting everything together, the full interaction term (24) thus becomes
$S_{\mathrm{int}}$ $=-\eta\int\!\mathcal{D}\Phi_{<}P[\Phi_{<}]e^{-\frac{1}{2}\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}\right)}$ (30)
×$\times\exp\left\{\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\left[\frac{1}{\sigma^{2}}\left(\sigma^{-2}+C^{-1}\right)^{-1}-\mathbb{1}\right]\Phi_{<}\right\}$
$=-\eta\,\,e^{-\frac{1}{2}\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}\right)}\!$
@@
×$\times\int\!\mathrm{d}\mu_{x}\,e^{-\frac{1}{2}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)^{\top}\left(\sigma^{2}+C\right)^{-1}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)}\leavevmode\nobreak\ ,$
@@
where on the second line we have reversed our measure changes above, [^6] Explicitly, $\int\!\mathcal{D}\Phi_{<}P[\Phi_{<}]=\int\!\prod_{k\leq\kappa}\mathrm{d}\varphi_{k}\,P[\varphi]=\int\!\prod_{k\leq\kappa}\mathrm{d}\varphi_{k}\int\!\mathrm{d}\mu_{x}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)=\int\!\mathrm{d}\mu_{x}$. and $\bm{f}_{m<}(x)$ and $\bm{y}_{<}(x)$ are $M$-dimensional vectors with elements $f_{m<}(x)$ and $y_{<}(x)$, respectively, cf. (13) and (14). Notably, in this expression all the higher feature modes $\phi_{k>\kappa}$ have been removed, and all higher GP modes $f_{m>}$ enter only through $C$, cf. (28). We note in passing that replacing $C$ with its self-consistent mean-value leads to the same average predictor as in [5].
@@
Next, we recall that, as the replica limit ($M\to 0$) is taken, our partition function should provide us with the average and variance of $f_{mk}$ as obtained by GP regression on fixed data, averaged over all datasets. Notably, the average and variance are respectively first and zeroth order in $y$ per dataset and hence must remain so after this averaging. However, in contrast to this expected result, the interaction term in our setting involves arbitrarily high powers of $y$. We argue in appendix B that the expected linear behavior in $y$ is restored in the interaction term in the replica limit. An easy way to see this is would be to start with the interaction term having a scaled $\epsilon\,y_{>}$ for an infinitesimal parameter $\epsilon$, followed by divisions by the same. Any term containing second or higher order expressions in $\epsilon\,y_{>}$ would vanish, leaving aside linear terms in $y_{>}$. Following these arguments, when calculating the average predictor (which is linear in $y_{<}$), we may set $y_{>}=0$ from the start, which we do henceforth. One can then show that the contribution $\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}_{M\times M}\right)$ vanishes in the replica limit. We therefore proceed with the following interaction term:
@@
We now wish to integrate out the higher GP modes still present in the covariance matrix. This is a complicated step due to nonlinear correlations between higher and lesser GP modes in (30). To proceed, we perform the Wilsonian RG procedure perturbatively as follows: we integrate out an infinitesimal momentum shell $\kappa-\delta\kappa\leq k<\kappa$, with the cutoff $\kappa$ sufficiently large such that it renders the sum of marginalized GP kernel eigenvalues $\lambda_{k}$ (cf. (7)) much smaller than the variance scale of the observation noise $\sigma^{2}$. That is, under the free theory for the higher modes, the expectation value of the GP covariance matrix (28) is [^7] To see this, observe from the first line of (15) that $\lambda_{k}$ is the variance of the Gaussian higher modes in $S_{0>}=\tfrac{1}{2}\sum_{k>\kappa}\lambda_{k}^{-1}f_{mk}^{2}$,viewed as a probability distribution in replica space. Since $[C_{y_{>}=0}]_{nm}=\sum_{k>\kappa}f_{mk}f_{nk}$, the expectation value $C_{y_{>}=0}\right>_{S_{0}}$ is the second cumulant of this distribution.
@@
This ensures that at each step of the RG, $[C_{y_{>}=0}]_{nm}=\sum_{k}f_{mk}f_{nk}$ has a typical scale of $\delta c$. Since $\delta c/\sigma^{2}\ll 1$, we may expand the interaction term, keeping only the leading-order contributions in $C_{y_{>}=0}$. This amounts to replacing $C_{y_{>}=0}$ in (31) by its average according to the free theory, namely $C_{y_{>}=0}\rightarrow\delta c\,\mathbb{1}_{M\times M}$. We thus find a renormalized ridge parameter $\sigma^{\prime}$, given by
@@
It is interesting to compare this result with other works which use the feature Gaussianity assumption, such as [5]. There it was found that the dataset average predictor coincides with the Equivalent Kernel approximation with an effective ridge parameter, $\kappa$. This is similar to our result in that their $\kappa$ is the original ridge parameter plus all the eigenfunctions which are unlearnable. There are two important conceptual differences however: in that derivation, there is no notion of RG, nor of a Wilsonian cut-off or RG flows. The same largely holds for their concurrent work [7]. Furthermore, the criteria for absorbing unlearnable modes is not sharp as in our case, but is instead controlled by the ratio of $\kappa$ over the number of datapoints and the eigenfunctions.
III.1 Neural Scaling Laws
@@
IV Renormalization beyond the Gaussian
@@
Here we extend the RG formalism to accommodate non-Gaussian effects associated with the feature modes, namely, cases where the $\varphi$ random variables (under the data measure) have substantial deviations from normality. We begin with a somewhat artificial but illustrative scenario where the greater features are non-Gaussian yet remain independent from the lesser features. In this case, we show that the previous treatment still holds without any changes. We then proceed with allowing weak non-Gaussianity with potential dependencies between greater and lesser feature modes. In this case, alongside the noise renormalization, we obtain a spatial re-weighting of the mean squared error (MSE) loss or, equivalently, a spatial re-weighting of the ridge parameter.
@@
In accordance with our perturbative RG prescription above, we take $\Phi_{>}$ to contain a single feature mode (the rest are to be integrated out in subsequent steps), which we denote by $\varphi_{q}$, so that $\Phi_{q}$ is the replica space vector whose $m^{\mathrm{th}}$ element is $\Phi_{mq}\coloneqq[\Phi_{q}]_{m}=(f_{mq}-y_{q})\varphi_{q}$, cf. (23). The above integral can then be viewed as the cumulant generating function of $P[\varphi_{q}]\exp\{-\frac{1}{2\sigma^{2}}\Phi_{q}^{\top}\Phi_{q}\!\}$ with $-\Phi_{q}^{\top}\Phi_{<}/\sigma^{2}$ playing the role of the cumulant function argument. Accordingly, the leading order non-Gaussian corrections will enter as $(f_{<}(x)-y_{<}(x))^{4}$ times the fourth cumulant of $\varphi_{q}$. For reasonable (e.g., sub-Gaussian) distributions, [^8] Since our approach is perturbative, we limit our study to distributions without heavy tails, cf. [38]. the (2n)th cumulant scales like the variance to the $n^{\mathrm{th}}$ power. The higher cumulants are thus $O\left(\delta c^{2}\right)$ and hence negligible, and we obtain the same equations as in the Gaussian case.
We now turn to the more realistic case in which the lesser and greater feature modes are coupled, and relax the Gaussianity assumption by noting that if the data has a high effective dimension, $d$, then a generic feature – which is a sum of many monomials in $x_{i}$ – can be viewed as a sum of many weakly correlated random variables and hence well-approximated by a Gaussian. This, together with the empirical success of the Gaussian assumption [5], motivates us to consider corrections to Gaussianity as small $1/d$ corrections (as in [39]) and treat the latter perturbatively. At a technical level, we do so by keeping a leading-order Edgeworth expansion of $P[\varphi_{<},\varphi_{>}]$. We then use this expansion to obtain the conditional distribution $P[\varphi_{>}|\varphi_{<}]$ appearing in (24), and perform the integral perturbatively by keeping only the leading non-Gaussianities. Concretely, we take as our large-$d$ expansion
$P[\varphi_{<},\varphi_{>}]\approx{\cal N}[0,\mathbb{1};\varphi_{<}]{\cal N}[0,\mathbb{1};\varphi_{>}]$
×$\times\Bigg{(}1+\frac{1}{4!}\sum_{\begin{subarray}{c}k_{1},k_{2},\\ k_{3},k_{4}\end{subarray}}U_{k_{1}k_{2}k_{3}k_{4}}He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})\Bigg{)}\leavevmode\nobreak\ ,\quad$ (37)
@@
$He_{4}$ $(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})$ (38)
$={\cal N}[0,\mathbb{1};\varphi_{<}]^{-1}{\cal N}[0,\mathbb{1};\varphi_{>}]^{-1}$
×$\quad\times\partial_{\varphi_{k_{1}}}\partial_{\varphi_{k_{2}}}\partial_{\varphi_{k_{3}}}\partial_{\varphi_{k_{4}}}{\cal N}[0,\mathbb{1};\varphi_{<}]{\cal N}[0,\mathbb{1};\varphi_{>}]\leavevmode\nobreak\ ,\quad$
(see for example [40]), and $U$ is the (fourth) cumulant/Ursell function,
@@
i.e., the connected 4-pt correlation function with respect to $P[\varphi]$. We emphasize that in our perturbative expansion in $1/d$, we shall keep only terms which are leading order in the fourth cumulant. Notably the lack of odd cumulants is consistent with our previous simplifying assumption that all features are odd functions therefore $P[\varphi]=P[-\varphi]$. In general, we think of $\varphi$ as spanning some relevant (i.e., potentially learnable) subspace of function space and do not require this subspace to be closed under the multiplication of features (e.g., $\phi_{k}(x)^{2}$, is an even function and thus outside this subspace). A different set of assumptions is taken in sec. (IV.2), leading to similar results.
@@
$P$ $[\varphi_{<},\varphi_{q}]$ (40)
$\approx{\cal N}[0,\mathbb{1};\varphi_{<}]\left(1+\frac{1}{4!}U_{k1k_{2}k_{3}k_{4}}He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})\right)$
@@
$He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}})$ $=\varphi_{k_{1}}\varphi_{k_{2}}\varphi_{k_{3}}-\delta_{k_{1},k_{2}}\varphi_{k_{3}}$ (42)
$\quad-\delta_{k_{2},k_{3}}\varphi_{k_{1}}-\delta_{k_{1},k_{3}}\varphi_{k_{2}}\leavevmode\nobreak\ ,$
@@
$P[\varphi_{q}|\varphi_{<}]$ $\equiv\frac{P[\varphi_{q},\varphi_{<}]}{P[\varphi_{<}]}$ (44)
$\approx{\cal N}[0,\mathbb{1};\varphi_{q}]\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right],$
@@
$A\leavevmode\nobreak$ $\coloneqq\frac{1}{3!}U_{k_{1}k_{2}k_{3}q}He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}}),$
$B\leavevmode\nobreak$ $\coloneqq\frac{1}{2!}U_{k_{1}k_{2}qq}He_{2}(\varphi_{k_{1}},\varphi_{k_{2}}),$ (45)

### v3

@@ III Renormalizing Gaussian features
$I_{0}[\Phi_{>}]+I_{\mathrm{int}}[\Phi_{<},\Phi_{>}]\coloneqq$ $\,\frac{1}{2\sigma^{2}}\sum_{m;kk^{\prime}>\kappa}(f_{mk}-y_{k})(f_{mk^{\prime}}-y_{k^{\prime}})\varphi_{k}\varphi_{k^{\prime}}$ (25)
@@
We can then express this result in terms of a covariance matrix of replica indices by Taylor expanding the logarithm in the coefficient, using the cyclicity of the trace, and re-summing the series to obtain
$\frac{1}{2}\mathrm{tr}\ln\left(\mathbb{1}_{N_{>}\times N_{>}}+F^{\top}F/\sigma^{2}\right)$ $=\frac{1}{2}\mathrm{tr}\left(\frac{FF^{T}}{2\sigma^{2}}-\frac{(FF^{\top})^{2}}{4\sigma^{2}}+\ldots\right)$ (26)
@@
where $C\coloneqq FF^{\top}$ is the desired $M\!\times\!M$ covariance matrix with elements
$[C]_{mn}=(f_{mk}-y_{k})(f_{nk}-y_{k})\leavevmode\nobreak\ ,$ (27)
@@
cf. the definition of $F^{\top}\!F$ in (25). The argument of the exponential can similarly be written in terms of this matrix by applying the Woodbury matrix identity, whereupon
$F\left(\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\right)^{-1}\!F^{\top}=\left(C^{-1}+\sigma^{-2}\right)^{-1}\leavevmode\nobreak\ ,$ (28)
@@
see appendix B. Putting everything together, the full interaction term (23) thus becomes
$S_{\mathrm{int}}=$ $-\eta\int\!\mathcal{D}\Phi_{<}P[\Phi_{<}]e^{-\frac{1}{2}\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}\right)}\exp\left\{\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\left[\frac{1}{\sigma^{2}}\left(\sigma^{-2}+C^{-1}\right)^{-1}-\mathbb{1}\right]\Phi_{<}\right\}$ (29)
@@
$=$ $-\eta\,\,e^{-\frac{1}{2}\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}\right)}\!\int\!\mathrm{d}\mu_{x}\,e^{-\frac{1}{2}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)^{\top}\left(\sigma^{2}+C\right)^{-1}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)}\leavevmode\nobreak\ ,$
@@
where on the second line we have reversed our measure changes above [^8] Explicitly, $\int\!\mathcal{D}\Phi_{<}P[\Phi_{<}]=\int\!\prod_{k\leq\kappa}\mathrm{d}\varphi_{k}\,P[\varphi]=\int\!\prod_{k\leq\kappa}\mathrm{d}\varphi_{k}\int\!\mathrm{d}\mu_{x}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)=\int\!\mathrm{d}\mu_{x}$., and $\bm{f}_{m<}(x)$ and $\bm{y}_{<}(x)$ are $M$-dimensional vectors with elements $f_{m<}(x)$ and $y_{<}(x)$, respectively, cf. (13) and (14). Notably, in this expression all the higher feature modes $\phi_{k>\kappa}$ have been removed, and all higher GP modes $f_{m>}$ enter only through $C$, cf. (27). We note in passing that replacing $C$ with its self-consistent mean-value leads to the same average predictor as in canatar .
@@
Next, we recall that, as the replica limit ($M\to 0$) is taken, our partition function should provide us with the average and variance of $f_{mk}$ as obtained by GP regression on fixed data, averaged over all datasets. Notably, the average and variance are respectively first and zeroth order in $y$ per dataset and hence must remain so after this averaging. However, in contrast to this expected result, the interaction term in our setting involves arbitrarily high powers of $y$. We argue in appendix C that the expected linear behavior in $y$ is restored in the interaction term in the replica limit. An easy way to see this is would be to start with the interaction term having a scaled $\epsilon\,y_{>}$ for an infinitesimal parameter $\epsilon$, followed by divisions by the same. Any term containing second or higher order expressions in $\epsilon\,y_{>}$ would vanish, leaving aside linear terms in $y_{>}$. Following these arguments, when calculating the average predictor (which is linear in $y_{<}$), we may set $y_{>}=0$ from the start, which we do henceforth. One can then show that the contribution $\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}_{M\times M}\right)$ vanishes in the replica limit. We therefore proceed with the following interaction term:
@@
As a mid-derivation summary, we have now integrated out the greater modes of the input measure leading to the above rewriting of the interaction term. However, we still need to integrate out the greater GP modes (i.e., $f_{mk}$ with $k$ above the cutoff) contained in the replica space matrix $(\sigma^{2}+C_{y_{>}=0})^{-1}$. Despite its seemingly complex appearance, the Wilsonian RG procedure allows us to carry out this integration perturbatively as follows: we integrate out an infinitesimal momentum shell $\kappa-\delta\kappa\leq k<\kappa$, with the cutoff $\kappa$ sufficiently large such that it renders the sum of marginalized GP kernel eigenvalues $\lambda_{k}$ (cf. (7)) much smaller than the variance scale of the observation noise $\sigma^{2}$. That is, under the free theory for the higher modes, the expectation value of the GP covariance matrix (27) is [^9] To see this, observe from the first line of (15) that $\lambda_{k}$ is the variance of the Gaussian higher modes in $S_{0>}=\tfrac{1}{2}\sum_{k>\kappa}\lambda_{k}^{-1}f_{mk}^{2}$,viewed as a probability distribution in replica space. Since $[C_{y_{>}=0}]_{nm}=\sum_{k>\kappa}f_{mk}f_{nk}$, the expectation value $C_{y_{>}=0}\right>_{S_{0}}$ is the second cumulant of this distribution.
@@
This ensures that at each step of the RG, $[C_{y_{>}=0}]_{nm}=\sum_{k}f_{mk}f_{nk}$ has a typical scale of $\delta c$. Since $\delta c/\sigma^{2}\ll 1$, we may expand the interaction term, keeping only the leading-order contributions in $C_{y_{>}=0}$. One can then integrate out the greater GP modes to leading order in $C_{y_{>}=0}$, which amounts to replacing $C_{y_{>}=0}$ in (30) by its average according to the free theory of the greater modes, namely $C_{y_{>}=0}\rightarrow\delta c\,\mathbb{1}_{M\times M}$. We thus find a renormalized ridge parameter $\sigma^{\prime}$, given by
@@
It is interesting to compare this result with other works which use the feature Gaussianity assumption, such as canatar . There it was found that the dataset average predictor coincides with the Equivalent Kernel approximation with an effective ridge parameter, $\kappa$. This is similar to our result in that their $\kappa$ is the original ridge parameter plus all the eigenfunctions which are unlearnable. There are two important conceptual differences however: in that derivation, there is no notion of RG, nor of a Wilsonian cut-off or RG flows. The same largely holds for their concurrent work atanasov2024scaling . Furthermore, the criteria for absorbing unlearnable modes is not sharp as in our case, but is instead controlled by the ratio of $\kappa$ over the number of datapoints and the eigenfunctions.
In the next section, we will show that the Gaussian $P[\varphi]$ assumption leads to reasonable predictions for two real-world datasets (MNIST and CIFAR10). In section V, we will study how the picture changes when $P[\varphi]$ is allowed to be slightly non-Gaussian. As an illustrative empirical example, we then show how this non-Gaussian treatment can capture a qualitative effect which would otherwise be missed (section V.2).
IV Neural Scaling Laws
@@
IV.1 Empirical Demonstration
In this section we further corroborate the claim in section IV that our RG method is capable of predicting the power-law scaling behavior of the loss as a function of dataset size, which has been observed in real-world datasets bahri2021explaining . We consider one of the regression tasks from Ref. canatar on two such datasets, MNIST MNIST and CIFAR10 CIFAR10 . In these experiments, we compare the empirical scaling of the MSE loss to the predictions from our RG theory [^10] Our RG theory prediction involves two steps: (1) we find the renormalized noise, $\sigma_{eff}^{2}$, arising from unlearnable modes; (2) we use this renormalized noise, as opposed to the typical $\sigma^{2}$, in the EK approximation formula (11). Thus, by RG theory, we mean the EK prediction modified with a renormalized noise, $\sigma^{2}\rightarrow\sigma_{eff}^{2}$., the naive EK approximation (in which the noise, $\sigma^{2}$, is not renormalized), and the current state-of-the-art Spectral Bias theory canatar . We find that our RG theory accurately predicts the empirical power-law scaling of the MSE loss on these real-world datasets.
[…]
$\hat{y}(x)=\sum_{k=1}^{N}\leavevmode\nobreak\ \phi_{k}(x)\leavevmode\nobreak\ \bar{f}_{k}=\sum_{k=1}^{N}\left(\frac{\lambda_{k}}{\lambda_{k}+\sigma^{2}_{\rm eff}/\eta}\right)\leavevmode\nobreak\ \phi_{k}(x)\leavevmode\nobreak\ y_{k}.$ (35)
Comparing with (11), we can see that the key difference from the naive EK is the noise renormalization i.e. $\sigma^{2}\rightarrow\sigma_{eff}^{2}$. We can then express the theoretical MSE loss as follows
${\rm MSE}\left(y,\hat{y}\right)=\frac{1}{N}\sum_{k=1}^{N}\lvert\leavevmode\nobreak\ \bar{f}_{k}-y_{k}\rvert^{2}=\frac{1}{N}\sum_{k=1}^{N}L_{k}^{2}\leavevmode\nobreak\ y_{k}^{2},$ (36)
where we have implicitly defined the learnability factor,
$L_{k}:=\frac{\sigma^{2}_{\rm eff}}{\eta\lambda_{k}+\sigma^{2}_{\rm eff}}.$ (37)
In section IV, we assumed that $L_{k}=0$ for a learnable mode ($k\leq\kappa$) and $L_{k}=1$ for an unlearnable mode ($k>\kappa$). Here we will offer a more careful treatment using (37). Fixing a learnability threshold, $T\in(0,1)$, we numerically find the number of learnable modes, $\kappa$, such that $L_{\kappa}\approx T$ [^11] In appendix E, we explore the sensitivity to different threshold choices.. Note that we assume $1\leq\kappa\leq N$ (i.e. that there is at least one learnable mode). Thus, $\kappa$ generally will depend both on the number of observations, $\eta$, as well as the initial noise, $\sigma^{2}$, which enters through $L_{k}$’s dependence on $\sigma_{\rm eff}^{2}=\sigma^{2}+\sum_{k=\kappa+1}^{N}\lambda_{k}$. With this critical $\kappa_{T}$, we then calculate our prediction for the loss according to (36). Repeating this procedure for increasing numbers of observations, $\eta$, gives us a theoretical prediction of how the MSE loss scales as a function of dataset size. In Fig. 1, we compare these theoretical predictions to empirical GP kernel regression results obtained using the scheme described in Ref. canatar . Indeed, we find that our RG theory accurately predicts the power-law dependence of the loss on $\eta$, and is a great improvement over the naive EK prediction (i.e., (11)).
Figure 1: Predicting empirical MSE loss scaling on MNIST and CIFAR10 regression tasks. For two real-world dataset examples (MNIST and CIFAR10), we show different theoretical predictions for the MSE loss as a function of the number of datapoints, $\eta$. For each dataset, we consider two different regression tasks. For MNIST we consider both ‘8-9’ and ‘0-1’, where $N=9,837$ and $N=10,564$, respectively. For CIFAR10 we consider both ‘ship-truck’ and ‘automobile-bird’, where $N=10,000$ in both cases. For all experiments, we fix $\sigma^{2}=10^{-8}$. We find that both the state-of-the-art Spectral Bias theory canatar and the RG theory introduced in this work predict the semi power-law behavior well. Whereas, the EK approximation in which there is no noise renormalization (i.e. $\sigma_{\rm eff}^{2}=\sigma^{2}$) fails to accurately predict this behavior.
Figure 2: Illustrative depiction that the feature modes of real-world datasets are often Gaussian distributed. For two real-world dataset examples (MNIST and CIFAR10), we select two feature modes and show their joint distribution along with the best Gaussian fit to each marginal. This is done for each learning task considered. Namely, ‘8-9’ and ‘0-1’ in the MNIST case and ‘ship-truck’ and ‘automobile-bird’ in the CIFAR10 case. As previously noted in Simon2021 , we can observe the interesting fact that the feature modes of real-world datasets are often (jointly) Gaussian distributed. Appendix F explores the extent to which this varies as a function of $k$.
V Renormalization beyond the Gaussian
@@
Here we extend the RG formalism to accommodate non-Gaussian effects associated with the feature modes, namely, cases where the $\varphi$ random variables (under the data measure) have substantial deviations from normality. We begin with a somewhat artificial but illustrative scenario where the greater features are non-Gaussian yet remain independent from the lesser features; that is, $P[\varphi_{>}|\varphi_{<}]=P[\varphi_{>}]$ with $P[\varphi_{>}]$ non-Gaussian. In this case, we show that the non-Gaussianity of $P[\varphi_{>}]$ has no effect, and thus the previous treatment still holds without any changes. We then proceed with allowing weak non-Gaussianity with potential dependencies between greater and lesser feature modes. Schematically, this corresponds to $P[\varphi_{<},\varphi_{>}]\approx{\cal N}[0,\mathbb{1};\varphi_{<}]{\cal N}[0,\mathbb{1};\varphi_{>}]+\text{corrections}\Rightarrow P[\varphi_{>}|\varphi_{<}]\approx{\cal N}[0,\mathbb{1};\varphi_{>}]+\text{corrections}$, where the corrections are to be made precise below. This corrected form of $P[\varphi_{>}|\varphi_{<}]$ is subsequently used in (23) where, alongside the noise renormalization, we obtain a spatial re-weighting of the mean squared error (MSE) loss or, equivalently, a spatial re-weighting of the ridge parameter.
@@
In accordance with our perturbative RG prescription above, we take $\Phi_{>}$ to contain a single feature mode (the rest are to be integrated out in subsequent steps), which we denote by $\varphi_{q}$, so that $\Phi_{q}$ is the replica space vector whose $m^{\mathrm{th}}$ element is $\Phi_{mq}\coloneqq[\Phi_{q}]_{m}=(f_{mq}-y_{q})\varphi_{q}$, cf. (22). The above integral can then be viewed as the cumulant generating function of $P[\varphi_{q}]\exp\{-\frac{1}{2\sigma^{2}}\Phi_{q}^{\top}\Phi_{q}\!\}$ with $-\Phi_{q}^{\top}\Phi_{<}/\sigma^{2}$ playing the role of the cumulant function argument. Accordingly, the leading order non-Gaussian corrections will enter as $(f_{<}(x)-y_{<}(x))^{4}$ times the fourth cumulant of $\varphi_{q}$. For reasonable (e.g., sub-Gaussian) distributions [^12] Since our approach is perturbative, we limit our study to distributions without heavy tails, cf. zavatoneveth2021exact ., the (2n)th cumulant scales like the variance to the $n^{\mathrm{th}}$ power. The higher cumulants are thus $O\left(\delta c^{2}\right)$ and hence negligible, and we obtain the same equations as in the Gaussian case.
We now turn to the more realistic case in which the lesser and greater feature modes are coupled, and relax the Gaussianity assumption by noting that if the data has a high effective dimension, $d$, then a generic feature – which is a sum of many monomials in $x_{i}$ – can be viewed as a sum of many weakly correlated random variables and hence well-approximated by a Gaussian. This, together with the empirical success of the Gaussian assumption canatar , motivates us to consider corrections to Gaussianity as small $1/d$ corrections (as in tomasini2022failure ) and treat the latter perturbatively, analogous to the 1/N expansion in quantum field theory. At a technical level, we do so by keeping a leading-order Edgeworth expansion of $P[\varphi_{<},\varphi_{>}]$. We then use this expansion to obtain the conditional distribution $P[\varphi_{>}|\varphi_{<}]$ appearing in (23), and perform the integral perturbatively by keeping only the leading non-Gaussianities. Concretely, we take as our large-$d$ expansion
$P[\varphi_{<},\varphi_{>}]\approx\;{\cal N}[0,\mathbb{1};\varphi_{<}]{\cal N}[0,\mathbb{1};\varphi_{>}]\Bigg{(}1+\frac{1}{4!}\sum_{\begin{subarray}{c}k_{1},k_{2},\\ k_{3},k_{4}\end{subarray}}U_{k_{1}k_{2}k_{3}k_{4}}He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})\Bigg{)}\leavevmode\nobreak\ ,\quad$ (39)
@@
$He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})=$ $\;{\cal N}[0,\mathbb{1};\varphi_{<}]^{-1}{\cal N}[0,\mathbb{1};\varphi_{>}]^{-1}\partial_{\varphi_{k_{1}}}\partial_{\varphi_{k_{2}}}\partial_{\varphi_{k_{3}}}\partial_{\varphi_{k_{4}}}{\cal N}[0,\mathbb{1};\varphi_{<}]{\cal N}[0,\mathbb{1};\varphi_{>}]\leavevmode\nobreak\ ,\quad$ (40)
(see for example Naveh_2021 ), and $U$ is the (fourth) cumulant/Ursell function,
@@
i.e., the connected 4-pt correlation function with respect to $P[\varphi]$  [^13] Note that this is the exact 4-pt function, which takes into account any potential RG effects in the input/$P[\varphi]$ distribution such as those considered in Ref. Bradde_2017 .. We emphasize that in our perturbative expansion in $1/d$, we shall keep only terms which are leading order in the fourth cumulant, again in analogy with the $1/N$ expansion, in which the connected $2k$-point correlation function is suppressed by $1/N^{k-1}$. Notably the lack of odd cumulants is consistent with our previous simplifying assumption that all features are odd functions therefore $P[\varphi]=P[-\varphi]$. In general, we think of $\varphi$ as spanning some relevant (i.e., potentially learnable) subspace of function space and do not require this subspace to be closed under the multiplication of features (e.g., $\phi_{k}(x)^{2}$, is an even function and thus outside this subspace). A different set of assumptions is taken in sec. (V.2), leading to similar results.
@@
$P[\varphi_{<},\varphi_{q}]\approx$ $\;{\cal N}[0,\mathbb{1};\varphi_{<}]\left(1+\frac{1}{4!}U_{k1k_{2}k_{3}k_{4}}He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})\right)$ (42)
@@
$He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}})$ $=\varphi_{k_{1}}\varphi_{k_{2}}\varphi_{k_{3}}-\delta_{k_{1},k_{2}}\varphi_{k_{3}}-\delta_{k_{2},k_{3}}\varphi_{k_{1}}-\delta_{k_{1},k_{3}}\varphi_{k_{2}}\leavevmode\nobreak\ ,$
@@
$P[\varphi_{q}|\varphi_{<}]\equiv\frac{P[\varphi_{q},\varphi_{<}]}{P[\varphi_{<}]}\approx{\cal N}[0,\mathbb{1};\varphi_{q}]\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right],$ (45)
@@
$A$ $\coloneqq\frac{1}{3!}U_{k_{1}k_{2}k_{3}q}He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}}),$ (46)
$B$ $\coloneqq\frac{1}{2!}U_{k_{1}k_{2}qq}He_{2}(\varphi_{k_{1}},\varphi_{k_{2}}),$